OpenStudy (anonymous):

Pre calculus help? I need to learn how to find the x-intercepts of logarithmic functions. Here is one problem I need to find it for: h(x) = log4(x-3)

OpenStudy (ksaimouli):

multiplyt by 4

OpenStudy (ksaimouli):

is 4 base

OpenStudy (anonymous):

Yeah, 4 is the base.

OpenStudy (anonymous):

plug in 0 for h(x) and solve for x so 0=log4(x-3) 4^0=x-3 (prperty of logs) so 1=x-3 x=4 x-itercept at x=4

OpenStudy (anonymous):

Thanks so much. :)

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Can I ask one more question?

OpenStudy (anonymous):

@jamesumner93

OpenStudy (anonymous):

yea sure